QUESTION 16

  1. Monica purchased 100 shares of NOLA mutual fund at a price of $25 per share at the beginning of the year and paid a front-end load of 4.5%. The securities in the fund increased in value by 11% during the year and the fund’s expense ratio was 1.2%. What was Monica’s annual return for this year assuming she were to sell at year end?

    a. 4.50%.

    b. 4.85%.

    c. 5.30%.

    d. 9.80%.

QUESTION 26

  1. Most actively managed mutual funds, when compared to a market index:

    a. Beat the index return in all years.

    b. Beat the index return in most years.

    c. Exceed the return of the index.

    d. Do not generally outperform the index.
Answer #1

Effective purchase price = (1 - 0.045)(25) = $23.875

Price of share after all expenses = 23.875(1.11)(1 - 0.012) = $26.18

Return Rate = (26.18 - 25)/25

Return Rate = 4.85%

Do not generally outperform the index